# A Real-time Multimodal Transformer Neural Network-powered Wildfire Forecasting System

### v1

@@ 3.4 Hybrid Multimodal Model
In the hybrid model, the imagery data will be processed by the previous model and image features and token features will be the outputs from the process. Then the outputs will concatenate with the information data and fed the new concatenated data into the baseline model. Thus, as we can see from Figure 13, the hybrid model will take the results from wildfire images information net, then together with the information data into the baseline model. At the last stage, the output of the model will be the required target probabilities. In the metrics, we set the threshold as $0.5$ which means if the probability is large than $0.5$ then we categorized as true, otherwise false. Mathematically,
@@
where $\hat{R}$ are the responses, $p$ is the outputs from the model. During the training, it usually reaches top $TPR$ and $TNR$ at about $100$ epochs. In the hybrid model, we totally have $268615$ variables.
@@ 4.1 Forecast Metrics
The metrics we care most and used to measure the performances is the true positive rate (TPR or sensitivity) and true negative rate (TNR or selectivity)[47], which both are the components of accuracy. True positive is defined as values that are actually positive and predicted positive. In this case the true positive value provides the results that given a location with wildfire, it is caused by natural, and the prediction is also caused by natural, thus, we would like this value as large as possible. On the contrary, false negative value provides the results that given a location occurred wildfire, model predict it is caused by natural while it is in fact not, which is a type II error, thus we would like to avoid. Therefore, with these two values, true positive rate can provide us a good metrics on the performance on for the model we would like this value as large as possible. Since without undersampling method, model tends to have a high true negative rate, thus we try to find an optimal values between true positive rate and true negative rate. It provides a conditional probability that if there is a wildfire, the model can categorized whether it is caused by nature or not. Mathematically, we have
@@
There are other metrics often used in analysis, F-score[48], which can be represented as
$F-score$ $=\frac{True\ Positives}{True\ Positives+0.5(False\ Positive+False\ Negative)}$
and precision or positive predictive value,
$Precision=\frac{TP}{TP+FP}$
However, we do not include F-score and precision since in our test case the false counts is much more than the true one as we can see from the equation, the F-score gives equal importance to precision and recall[49]. The precision is related to the proportion of true cases and false cases[50], (see 6. Precision), in our case, since the precision will lead some waste of resources, our top priorities are still $TPR$ and $TNR$. Using the terminology coined in this paper, as can be seen from Eq. (5), a high “True Positive” rate means that when the model predicts a fire in a certain place, there is actually a fire occurs, while more false negative indicates a low rate of “True Positive”, which means the possibility of mis-classifying the real occurred wildfire. The term “False Positive” in our terminology means that there is no fire, and the model correctly categorizes it as no fire occurrence. If this rate is low, it indicates that there will be some false alerts as shown in Eq. (6), which may waste resources while ignoring the alert might be detrimental. Thus, currently our focus will be first increase the $TPR$, the $TNR$ and hence the accuracy in reliability.
@@ 4.2 Baseline Model Results
The baseline model with recommended setting (See: 8) provides a relatively stable results for all the baseline cases we trained. As we can see from Figure 14 (a), Figure 15 (a), and Figure 16 (a), 5 pre-trained results are chosen which are also available in Colab. The variance of the losses of each cases are relatively low for 5 random runs. However, with different selected features, the results are slightly improved. As one can see from Figure 14 (a), the losses are successfully converged to a small but little higher values than $0.06$. On the accuracy side, as shown in Figure 14 (b), the test accuracy at the last epoch 300 is all within the range of $89\%$ to $90\%$.
@@
If we add vegetation categories as new features to training dataset, with the same architecture with the hyperparameters, the results will be slightly improved as shown in Figure 15 (a). Compare with the previous one, the loss is slightly smaller, which mostly are within 0.06, as shown in Figure 17, the line in red is the results for not include vegetation category and the line in orange is the one with vegetation category. Same situation occurs in test accuracy as shown in Figure 15 (b), we can see that the overall accuracy is better and some of the results are more than $90\%$ accuracy. Model Stability is one of the core principles in data science [51]. For each of the models, we randomly split train and test data, as shown in all errors, the baseline model is relatively stable. We can also see from Figure 17 (b) that the accuracy is even better if the features including the vegetation categories.
@@
As we can see from Figure 16 (a), if we use undersampling data for training the losses are higher than previous case, and the accuracies are also lower, which is around $82\%$. The reason for the high accuracy is due to the imbalance of data, the model tends to categorize as false, since more than $80\%$ data are false, thus, model fits well for categorizing most case as false.
@@
As shown in Figure 18 (a), we can see that among $5910$ test data, $5015$ are categorized as negative which leads a $94.6\%$ TNR. Even though a little better with vegetation categories, similar situation occurs in the cases with vegetation category, as shown in Figure 18 (b), among the same amount test data, $5910$, $4985$ are categorized as false. For sensitivity, the results are similar to that of without vegetation category, which is also more than $90\%$. For specificity, compare to previous one, we can see that the results have about $5\%$ better. Even with limited improvement, this results are better than the previous. Thus specificity contributed most accuracy, since it has large proportion and relatively high results. As we mentioned, sensitivity is the top priority and the part that we care most, by using these features, the sensitivity is not high enough even though it provided a high accuracy. Thus if we only using with 7 days, 15 days, and 30 days average temperature and wind data as features, the neural network model converge but with around $89\%$ accuracy and quite low sensitivities ($65\%-70\%$) with a high specificity with more than $90\%$, as shown in Figure 18 (a). After applying undersampling method, we can see the results from Figure 18 (c), even though the weighted accuracy is lower than previous cases, the $TPR$ increases a lot to more than $80\%$. For total $18985$ test cases, model found an optimal solution to balance the $TPR$ and $TNR$ which are all around $80\%$ which leads the total accuracy also about $80\%$. This is the desired results we would like to see.
@@
Then if we adding 7 days, 15 days, and 30 days average humidity or precipitation data, the neural network model converge and losses are around 0.07 and accuracy is around $88\%-89\%$, the confusion matrix is not good since the sensitivities are less than $60\%$. Once we adding vegetation type data, the losses are decreased about $0.01$ to $0.05-0.06$ and accuracy will increase around $1\%$, the sensitivities of confusion matrix increases to around more $60\%-65\%$.
@@ 4.3 Hybrid model results
Figure 19 shows the how some images change during different stages from a pre-trained model, we can see the top of Figure 19 is the original grey-scaled image and the output of the first convolutional layer of the same index are the middle images. After the first trained convolutional layer, the trained results successfully maintain the features from the original images. We are able to see the obvious features from the previous original images. While for the bottom of Figure 19, due to the resolution is quite small, which is $10\times 10$, we can see some similarities if compare with the original images. Thus, when finishing training, the convolutional part do extracting feature maps from the images and feed them to next stage.
@@
Figure 20 is the output of pre-trained model with same index from Figure 19. The original images has a resolution of $5\times 5$, even with a small resolution, we believe that this is how the model sees the images and finds the features of a certain condition for the model. Figure 7. Example of Sample Outputs shows detailed outputs for each layers.
@@
As we can see from Figure 21, the loss successfully converged and reached in a range of $0.07$ to $0.08$. Our top priority, $TPR$ reaches $85\%$ while the $TNR$ also reached $85\%$ during that within that range of losses. Therefore, our accuracy is also around $85\%$ as calculated by $\frac{TPR+TNR}{P+N}$, as shown in Figure 22. However, we can see from both of the confusion matrix in Figure 18 (c) and Figure 22, the precisions are relatively low since inside the test case for undersampling, we move some majorities to test case. This is one of the drawbacks of this model. If we have a low precision it means that it has a high probability such that there would not be a fire, but we categorized as true that causes a false alarm and type II error will occurred.
@@
Usually, based on the experiences, during training, about $60$ epochs, the $TPR$ and $TNR$ will both reach around $80\%$, so does accuracy. When reaching $100$ epochs, all of them might be more than $83\%$.
@@ 4.4 Main results
From Table 5, we summarized the results of our experiments. As shown, the best $TNR$ is in all the features without images and undersampling method. Even though With a $TPR$ about $70\%$, the precision is relative high, which means that the model can predict accurately within the true cases. However, there is a trade-off, with the $TPR$ increase, our $TNR$ experiment results go down $10\%$, as shown in Table 5. So does the precision, since the precision is related to the proportion of our data as shown in 6. Precision, among the $18955$ test cases, there are only $5\%$ are with the true label, even $85\%$ of them are categorized right, the precision, $\mathbb{P}(actual=True|predict=True)$, went down to about $23.86\%$ from Figure 22, with a $5\%$ increase from the results, $18.78\%$, of the features without images.
@@
As we can see from Figure 23 (b), another optimal case, the results without add on the majorities to the test case we still have about $85\%$ accuracy. Since all true cases are in the original test cases the $TPR$ will not change. We can see that from both Figures 23 (a) and 23 (b), for the added majorities, which has $16343$ cases with all false, the TN is about $13885$ and FP is $2458$, Thus, the $TNR$ is about $84.96\%$ which is also the accuracy in this case since there are no true results in this part. However, if we do not include the majority test cases, the precision can reach about $\frac{787}{245+787}=76.26\%$, since our overall accuracy is still about $85\%$ with the more false cases the precision will go down. Therefore, the proportion of the data will has a large influences on the precision. If we include more true data late, the precision should go up. Moreover, once both $TPR$ and $TNR$ are improved, the precision should also be improved.
@@ 4.5 Case Study: 2018 California Camp Fire
To demonstrate and validate the ability and effectiveness of the proposed high-resolution multimodal transformer neural network model to forecast or predict the wildfire occurrence, we use the local weather forecasting data on November 7th, 2018, and the local Google Earth image data before November 7th, 2018, in the vicinity of Pulga, California, just right before the occurrence of the 2018 California Camp (November 8th, 2018) to test the performance of the proposed Bayesian machine learning model. Figure 24 is a real satellite image taken right after the starting occurrence of the California Camp fire near the town Pugla, California, on November, 8th 2018 by NASA. The two fire staring or occurrence locations are marked as colored circles (I: Orange circle and II: Red circle.) in Figure 24. By inputting the local weather data and tje image data from the Google Earth into the multimodal transform neural network, the multimodal transformer machine learning neural network will output the probability distribution in the local region. Figure 25 is the prediction result obtained by using the proposed multimodal transformer neural network model on prediction of the occurrence of the California Camp fire at the observed fire occurrence site I. The topographic and weather information input data and parameters are given as follows,
@@
where the first two are the converted longitude and latitude of the location. The next 12 parameters are the average temperature, wind speed, and precipitation, humidity for prior 7, 15, and 30 days before November, 8th 2018. The last 6 are the vegetation categories for this location. This row is repeated 195 times since this area share the same data value, the only difference is the 195 different input images as shown in Figure 25.
@@
The predicted wildfire probabilities on each small-scale box of $100m\times 100m$ are depicted in Figure 25. We marked all the wildfire occurrence high probability ($>70\%$) regions with yellow color probability numbers. For the fire occurrence location II, we also input the local geographic and weather data, as well as the Google Earth image data into the Bayesian machine learning neural network. The input information data is given as follows, [ 39.825097,-121.446211, 18.74354839, 18.240000, 18.30714286,
@@
The predicted wildfire probabilities on each small-scale box of $100m\times 100m$ are depicted in Figure 26. Again, we marked all the wildfire occurrence high probability ($>70\%$) regions with yellow color probability numbers as shown in Figure 26. It is worthy noting that the training data that we used is from 1992 to 2015, while the California Camp Fire occurred in 2018. Therefore, this example is a bona fide forecasting demonstration.
@@
From Figures 25 and 26, one may observe that all the places where a wildfire most likely occur, are regions where have dense vegetation and flat terrain conditions.
@@ 5 Conclusions and Perspectives
In this work, we developed a high-resolution multimodal transformer neural network model to forecast wildfire hazard occurrence location and timing at small scale. Different from the existing ML models for wildfire prediction, the present approach combines weather information, historical fire data, local topographic information, and Google Earth image information to train the multimodal transformer neural network machine learning model, and then practically provides the wildfire occurrence forecast at a local spatial location within $\leq 100m^{2}$. By using the wildfire data in the United States from 1992 to 2015 (USDA[9]) to train the high-resolution multimodal transformer neural network, we are able to provide the small scale spatial probability of wildfire occurrence with the input of the given weather forecasting information and the synchronized local Google Earth imagery information 24 hours before. This is a significant advance in wildfire forecasting and prediction technology, and we have made a major step forward in wildfire occurrence forecasting. In this work, we use the wildfire data collected in the $FPA\_FOD$ database to demonstrate the viability and the accuracy of the proposed artificial intelligence-based wildfire forecast system. In the fourth edition of $FPA\_FOD$ ([7]), it contains 1.88 millions of entries of wildfire records from 1992 to 2015. Recently, a new version with 2.3 millions of entries of wildfire records are released to the public [53]. It contains a spatial database of the wildfires that occurred in the US from 1992 to 2022, which is the 7th edition of wildfire database to support the National Fire Program Analysis (FPA) system[53]. Thus, based on this new database, we can use a database with more data information, and more data features, and better quality data to train the proposed Bayesian machine learning neural network. The new version database increases about $42$ thousand wildfire records within $5$ years. As mentioned before, some of the images from Google Earth are in gray scale before $2005$, however, with this newly released data, we should be able to use the colored images of vegetation and extract terrain topological feature. Moreover, all the occurrences from $FPA\_FOD$ are within the US, and recently, a data called Wildfire Prediction Dataset (Satellite Images) was also released[54]. It is based on wildfire data from Canada’s Open Government Portal[55]. It contains two categories, wildfire images and no wildfire images, we can used the wildfire images to increase and enrich data-set. Furthermore, we may even be able to add more data such as Global Fire Emissions Database version 4[56]. Combining with and extending remote sensing data[57] that collect from MODIS (Moderate Resolution Imaging Spectroradiometer). Besides the data collected from North America. the wildfire data in may also be available[58]. With more data, not only the neural network model could perform better, the transformer like model will also have a better performance since as we mentioned before, a transformer usually requires a large amount of data[41]. As mentioned in 2.4, the accuracy of the data may significantly affect the prediction outcome. Therefore, for a certain location if the data qualities could be better, especially for the image data, the accuracy could significantly increased. In the present work, some of the features that are still missing from the database, such as elevation or sensing data. Remote sensing data can also be added to the data if the amount data collected is enough, which can also help predicting wildfire occurrence[59]. We shall report an updated version of the multimodal transformer neural network model for wildfire prediction in a subsequent work.

### v2

@@ 3.4 Hybrid Multimodal Model
In the hybrid model, the imagery data will be processed by the previous model and image features and token features will be the outputs from the process. Then the outputs will concatenate with the information data and fed the new concatenated data into the baseline model. Thus, as shown from Figure 13, the hybrid model will take the results from wildfire images information net, then together with the information data into the baseline model. At the last stage, the output of the model will be the required target probabilities. In the metrics, the threshold is $0.5$, which means if the probability is larger than $0.5$, then the model is categorized as true or false. Mathematically, we can write
@@
where $\hat{R}$ are the responses, $p$ is the outputs from the model. During the training, it usually reaches the top $TPR$ and $TNR$ at about $100$ epochs. In the hybrid model, the model totally has $268615$ variables.
@@ 4.1 Forecast Metrics
The metric used to measure the performance is the True Positive Rate (TPR or sensitivity) and true negative rate (TNR or selectivity) [47], which both are the components of accuracy. True positive is defined as values that are positive and predicted positive. In this case, the True Positive value indicates a scenario in which a wildfire that occurred in a given location was caused by natural causes, and the prediction also indicates that the wildfire is caused by natural causes. Thus, this value should be as large as possible. On the contrary, a False Negative value represents a scenario in which a wildfire occurred in a given location, and the model predicts that it is due to natural causes, while they are not. This is a type II error; thus, it should be avoided. Therefore, a true positive rate with these two values can provide us with good performance metrics, Since the model tends to have a high true negative rate without using the undersampling method, finding optimal values between the true positive rate and the true negative rate is a top priority. It provides a conditional probability that if there is a wildfire, the model can categorize whether it is caused by nature or not. The metric used to measure the performance is the True Positive Rate (TPR or sensitivity) and True Negative Rate (TNR or selectivity) [47], which both are the components of accuracy. True positive is defined as values that are positive and predicted positive. In this case, the True Positive value indicates a scenario in which natural causes caused a wildfire in a given location, and the model’s prediction also indicates that natural causes caused the wildfire. Thus, this value should be as significant as possible. On the contrary, a False Negative value represents a scenario in which a wildfire occurred in a given location, and the model predicts that it is due to natural causes, while they are not. This is a type II error; thus, it should be avoided. Therefore, a True Positive Rate with these two values can give us good performance metrics. Since the model tends to have a high true negative rate without the undersampling method, finding optimal values between the True Positive Rate and the True Negative Rate is a top priority. It provides a conditional probability that if there is a wildfire, the model can categorize whether it is caused by nature. Mathematically, we have
@@
There are other metrics often used in the analysis, F-score [48], which can be represented as
${\rm F{-score}}$ $=\frac{True\ Positives}{True\ Positives+0.5(False\ Positive+False\ Negative)}$
and the precision or the positive predictive value, i.e.
$Precision=\frac{TP}{TP+FP}~{}.$
However, F-score and precision are not included here because, in our test case, the false counts are much more than the true ones, as shown in the equation. The F-score gives equal importance to precision and recall [49]. The precision is related to the proportion of true cases and false cases [50], (See Appendix 6. Precision). In our case, since precision will lead to some waste of resources, our top priorities are still $TPR$ and $TNR$. Using the terminology coined in this paper, as can be seen from Eq. (5), a high “True Positive” rate means that when the model predicts a fire in a certain area, there is actually a fire occurs, while the false negative indicates a low rate of “True Positive”, which means the possibility of misclassifying the real occurred wildfire. The term “False Positive” in our terminology, means that there is no fire, and the model correctly categorizes it as no fire occurrence. If this rate is low, it indicates that there will be some false alerts as shown in Eq. (6), which may waste resources while ignoring the alert might be detrimental. Thus, currently, our focus will be first to increase the $TPR$, the $TNR$ and hence the accuracy in reliability.
@@ 4.2 Baseline Model Results
The baseline model with recommended setting (See: 8) provides relatively stable results for all the baseline cases we trained. As we can see from Figure 14 (a), Figure 15 (a), and Figure 16 (a), 5 pre-trained results are chosen, which are also available in Colab. The variance of the losses of each case is relatively low for 5 random runs. However, with different selected features, the results are slightly improved. As one can see from Figure 14 (a), the losses are successfully converged to a small but little higher value than $0.06$. On the accuracy side, as shown in Figure 14 (b), the test accuracy at the last epoch 300 is all within the range of $89\%$ to $90\%$.
@@
If we add vegetation categories as new features to the training dataset, with the same architecture as the hyperparameters, the results will be slightly improved as shown in Figure 15 (a). Compared with the previous one, the loss is slightly smaller, mostly within 0.06, as shown in Figure 17, the line in red is the result of not including the vegetation category, and the line in orange is the one with the vegetation category. The same situation occurs in test accuracy as shown in Figure 15 (b), we can see that the overall accuracy is better and some of the results are more than $90\%$ accuracy. Model Stability is one of the core principles in data science [51]. For each of the models, we randomly split train and test data, as shown in all errors, the baseline model is relatively stable. We can also see from Figure 17 (b) that the accuracy is even better if the features include the vegetation categories.
@@
As we can see from Figure 16 (a) if we use the undersampling data for training, the losses are higher than previous case, and the accuracies are also lower, which is around $82\%$. The reason for the high accuracy is due to the imbalance of data. The model tends to categorize as false since more than $80\%$ data are false. Thus, the model fits well for categorizing most cases as false.
@@
As shown in Figure 18 (a), we can see that among $5910$ test data, $5015$ are categorized as negative, which leads to a $94.6\%$ TNR. Even though it is a little better with vegetation categories, a similar situation occurs in the cases with vegetation category, as shown in Figure 18 (b), among the same amount of test data, $5910$ and $4985$ are categorized as false. For sensitivity, the results are similar to those without the vegetation category, which is also more than $90\%$. For specificity, compared to the previous one, we can see that the results are about $5\%$ better. Even with limited improvement, these results are better than the previous. Thus, specificity contributed the most accurately since it has a large proportion and relatively high results. As we mentioned, sensitivity is the top priority and the part that we care about most, by using these features, the sensitivity is not high enough even though it provides a high accuracy. Thus, if we only used 7 days, 15 days, and 30 days’ average temperature and wind data as features, the neural network model converges but with around $89\%$ accuracy and quite low sensitivities ($65\%-70\%$) with a high specificity with more than $90\%$, as shown in Figure 18 (a). After applying the undersampling method, we can see the results from Figure 18 (c), even though the weighted accuracy is lower than in previous cases, the $TPR$ increases a lot to more than $80\%$. For a total of $18985$ test cases, the model found an optimal solution to balance the $TPR$ and $TNR$, which are around $80\%$, leading to the total accuracy also about $80\%$. This is the desired result.
@@
Then if we add 7 days, 15 days, and 30 days of average humidity or precipitation data, the neural network model converges, and losses are around 0.07 and accuracy is around $88\%-89\%$, the confusion matrix is not good since the sensitivities are less than $60\%$. Once we added vegetation-type data, the losses decreased by about $0.01$ to $0.05-0.06$ and accuracy will increase around $1\%$, the sensitivity of confusion matrix increases to around more $60\%-65\%$.
@@ 4.3 Hybrid model results
Figure 19 shows how some images change during different stages from a pre-trained model, we can see the top of Figure 19 is the original grey-scale image and the output of the first convolutional layer of the same index are the middle images. After the first trained convolutional layer, the trained results successfully maintain the features from the original images. We are able to see the obvious features from the previous original images. For the bottom of Figure 19, due to the resolution being quite small, which is $10\times 10$, we can see some similarities if compared with the original images. Thus, when finishing training, the convolutional part extracts feature maps from the images and feeds them to the next stage.
@@
Figure 20 is the output of the pre-trained model with the same index from Figure 19. The original images have a resolution of $5\times 5$, even with a small resolution, we believe that this is how the model sees the images and finds the features of a certain condition for the model. Figure 7. Example of Sample Outputs shows detailed outputs for each layer.
@@
As we can see from Figure 21, the loss successfully converged and reached in a range of $0.07$ to $0.08$. Our top priority, $TPR$ reaches $85\%$ while the $TNR$ also reached $85\%$ during that within that range of losses. Therefore, our accuracy is also around $85\%$ as calculated by $\frac{TPR+TNR}{P+N}$, as shown in Figure 22. However, we can see from both of the confusion matrices in Figure 18 (c) and Figure 22, the precisions are relatively low since inside the test case for undersampling, we move some majorities to the test case. This is one of the drawbacks of this model. If we have a low precision, it means that there is a high probability that there would not be a fire, but we categorized it as true, which causes a false alarm, and a type II error will occur.
@@
Usually, based on the experiences, during training, about $60$ epoch, the $TPR$ and $TNR$ will both reach around $80\%$ and so does accuracy. When reaching $100$ epochs, all of them might be more than $83\%$.
@@ 4.4 Main results
From Table 5, we summarized the results of our experiments. As shown, the best $TNR$ is in all the features without images and undersampling method. Even though with a $TPR$ about $70\%$, the precision is relatively high, which means that the model can predict accurately within the true cases. However, there is a trade-off, with the $TPR$ increase, our $TNR$ experiment results go down $10\%$, as shown in Table 5. So does the precision, since the precision is related to the proportion of our data as shown in Appendix 6. Precision, among the $18955$ test cases, there are only $5\%$ with the true label, even $85\%$ of them are categorized right, the precision, $\mathbb{P}(actual=True|predict=True)$, went down to about $23.86\%$ from Figure 22, with a $5\%$ increase from the results, $18.78\%$, of the features without images.
@@
As we can see from Figure 23 (b), another optimal case, the results without adding the majority to the test case, we still have about $85\%$ accuracy. Since all true cases are in the original test cases the $TPR$ will not change. We can see that from both Figures 23 (a) and 23 (b), for the added majorities, which has $16343$ cases with all false, the TN is about $13885$ and FP is $2458$, Thus, the $TNR$ is about $84.96\%$ which is also the accuracy in this case since there are no true results in this part. However, if we do not include the majority of test cases, the precision can reach about $\frac{787}{245+787}=76.26\%$, since our overall accuracy is still about $85\%$, with more false cases, the precision will go down. Therefore, the proportion of the data will have a large influence on the precision. If we include more true data later, the precision should go up. Moreover, once both $TPR$ and $TNR$ are improved, the precision should also be improved.
@@ 4.5 Case Study: 2018 California Camp Fire
To demonstrate and validate the ability and effectiveness of the proposed high-resolution multimodal transformer neural network model to forecast or predict wildfire occurrence, we used the local weather forecasting data on November 7th, 2018, and the local Google Earth image data before November 7th, 2018, in the vicinity of Pulga, California, just right before the occurrence of the 2018 California Camp (November 8th, 2018) to test the performance of the proposed Bayesian machine learning model. Figure 24 is a real satellite image taken right after the starting occurrence of the California Camp Fire near the town of Pugla, California, on November 8th, 2018, by NASA. The two fire staring or occurrence locations are marked as colored circles (I: Orange circle and II: Red circle.) in Figure 24. By inputting the local weather data and the image data from Google Earth into the multimodal transform neural network, the multimodal transformer machine learning neural network will output the probability distribution in the local region. Figure 25 is the prediction result obtained by using the proposed multimodal transformer neural network model on prediction of the occurrence of the California Camp Fire at the observed fire occurrence site I. The topographic and weather information input data and parameters are given as follows,
@@
where the first two are the converted longitude and latitude of the location. The next 12 parameters are the average temperature, wind speed, and precipitation, humidity for the prior 7, 15, and 30 days before November 8th, 2018. The last 6 are the vegetation categories for this location. This row is repeated 195 times since this area shares the same data value, the only difference is the 195 different input images as shown in Figure 25.
@@
The predicted wildfire probabilities on each small-scale box of $100m\times 100m$ are depicted in Figure 25. We marked all the wildfire occurrences with high probability ($>70\%$) regions with yellow color probability numbers. For the fire occurrence location II, we also input the local geographic and weather data, as well as Google Earth image data into the Bayesian machine learning neural network. The input information data is given as follows, [ 39.825097,-121.446211, 18.74354839, 18.240000, 18.30714286,
@@
The predicted wildfire probabilities on each small-scale box of $100m\times 100m$ are depicted in Figure 26. Again, we marked all the wildfire occurrences with high probability ($>70\%$) regions with yellow color probability numbers as shown in Figure 26. It is worth noting that the training data that we used is from 1992 to 2015, while the California Camp Fire occurred in 2018. Therefore, this example is a bona fide forecasting demonstration.
@@
From Figures 25 and 26, One may find that the places where wildfires most likely occur are regions with dense vegetation and flat terrain.
@@ 5 Conclusions and Perspectives
In this work, we developed a high-resolution multimodal transformer neural network model to forecast wildfire hazard occurrence location and timing at a small scale. Different from the existing ML models for wildfire prediction, the present approach combines weather information, historical fire data, local topographic information, and Google Earth image information to train the multimodal transformer neural network machine learning model, and then practically provides the wildfire occurrence forecast at a local spatial location within $\leq 100m^{2}$. By using the wildfire data in the United States from 1992 to 2015 (USDA[9]) to train the high-resolution multimodal transformer neural network, we can provide the small-scale spatial probability of wildfire occurrence with the input of the given weather forecasting information and the synchronized local Google Earth imagery information 24 hours before. This is a significant advance in wildfire forecasting and prediction technology, and we have made a major step forward in wildfire occurrence forecasting. Moreover, in this work, we use the wildfire data collected in the $FPA\_FOD$ database to demonstrate the viability and the accuracy of the proposed artificial intelligence-based wildfire forecast system. The fourth edition of $FPA\_FOD$ ([7]) contains 1.88 million entries of wildfire records from 1992 to 2015. Recently, a new version with 2.3 million entries of wildfire records was released to the public [53]. It contains a spatial database of wildfires that occurred in the US from 1992 to 2022, which is the 7th edition of the wildfire database to support the National Fire Program Analysis (FPA) system[53]. Thus, based on this new database, we can use a database with more data information, and more data features, and better-quality data to train the proposed Bayesian machine learning neural network. The new database version increases about $42$ thousand wildfire records within $5$ years. As mentioned before, some of the images from Google Earth are in grayscale before $2005$, however, with this newly released data, we should be able to use the colored images of vegetation and extract terrain topological features. Moreover, all the occurrences from $FPA\_FOD$ are within the US, and recently, a data called Wildfire Prediction Dataset (Satellite Images) was also released[54]. It is based on wildfire data from Canada’s Open Government Portal[55]. It contains two categories: wildfire images and no wildfire images. We can use wildfire images to increase and enrich the data set. Furthermore, we may even be able to add more data, such as Global Fire Emissions Database version 4[56]. Combining with and extending remote sensing data[57] that was collected from MODIS (Moderate Resolution Imaging Spectroradiometer). Besides the data collected from North America. Moreover, all the occurrences from $FPA\_FOD$ are within the US, and recently, a data called Wildfire Prediction Dataset (Satellite Images) was also released[54]. It is based on wildfire data from Canada’s Open Government Portal[55]. It contains two categories: wildfire images and no wildfire images. We can use wildfire images to increase and enrich the data set. Furthermore, we may even be able to add more data, such as Global Fire Emissions Database version 4[56]. Combining with and extending remote sensing data[57] that was collected from MODIS (Moderate Resolution Imaging Spectroradiometer). Besides the data collected from North America. the wildfire data in other parts of the world may also be available [58]. With more data, not only the neural network model could perform better, but the transformer like the model, will also have a better performance since, as we mentioned before, a transformer usually requires a large amount of data[41]. As mentioned in 2.4, the accuracy of the data may significantly affect the prediction outcome. Therefore, for a certain location, if the data quality could be better, especially for the image data, the accuracy could significantly be increased. In the present work, some of the features that are still missing from the database, such as elevation or sensing data. Remote sensing data can also be added to the data if the amount data collected is enough, which can also help predict wildfire occurrence[59]. We shall report an updated version of the multimodal transformer neural network model for wildfire prediction in subsequent work.